OpenStudy (anonymous):

how do i simplify 5 square root 320

OpenStudy (anonymous):

\[5\sqrt{320} = 40\sqrt{5}\]

OpenStudy (anonymous):

or \[\sqrt[5]{320} = \sqrt[5]{2^5*10} = 2\sqrt[5]{10}\]

OpenStudy (anonymous):

320 2 160 2 80 2 40 2 20 2 10 2 5 =(2^6)*5....5*2^3 = 5*8 = \[40\sqrt{5}\]